Grade 12th passPhysical Chemistry

Half life for zero order reaction , A(g)--------->B(g) + C(g) & half life for first order reaction X(g)-->Y(g) are equal. If completion time for zero order rxn is 13.86 min, then calculate rate constant for rxn. X----->Y

Arpit Dhankar
11 Years agoGrade 12th pass

1 Answer

Gopesh Kumar
11 Years ago
half life of the 1^{st} order reaction = 13.86/2 mins therefore the rate = 0.69/T^{_{1/2}} = 0.099 /min